# Revisiting Data Challenges of Computational Pathology: A Pack-based Multiple Instance Learning Framework

### v1

@@ B.6 Inference Time Comparison
B.7 Supplementary Confidence Intervals
@@
Table 17: Detailed CI for PackMIL-enhanced methods on all downstream tasks.
Method Benchmark CHIEF UNI GIGAP
@@
[…]
@@ Appendix C Qualitative Analysis
Figure 5: Attention visualization on the PANDA dataset Bulten et al. (2022). The RS strategy, due to its sampling, exhibits limited global attention. With multi-slide supervision via hyperslides and the supplementation of key features, PackMIL demonstrates a more accuracy and comprehensive focus on tumor areas.
@@ D.1 Task-specific Hyperslide Loss
Grading task. This task is modeled as single-label classification, we employ Asymmetric Loss (ASLLoss) Ridnik et al. (2021). This loss function is chosen to effectively address potential class imbalance and encourage the model to focus on correctly identifying positive classes while being less sensitive to negative misclassifications. Given the predicted probability distribution $\mathbf{p}=[p_{1},\dots,p_{G}]$ for $G$ classes and the corresponding one-hot encoded ground truth labels $\mathbf{y}^{\mathrm{hyper}}$, the loss is computed as:
$L_{\mathrm{grade}}(\mathbf{p},\mathbf{y}^{\mathrm{hyper}})=-\sum_{i=1}^{G}\mathbf{y}^{\mathrm{hyper}}_{i}(1-p_{i})^{\gamma_{p}}\log(p_{i})-\sum_{i=1}^{G}(1-\mathbf{y}^{\mathrm{hyper}}_{i})p_{i}^{\gamma_{n}}\log(1-p_{i}),$ (6)
@@
Subtyping task. This task is modeled as multi-label classification with soft targets, we use multi-label Focal Loss Lin et al. (2017). This loss helps mitigate the issue of imbalanced subtype frequencies and focuses the model’s attention on harder-to-classify samples. Given the predicted probability vector $\mathbf{p}=[p_{1},\dots,p_{C}]^{\top}$ for $C$ subtypes (typically obtained via Sigmoid activation) and the corresponding soft ground truth label vector $\mathbf{y}^{\mathrm{hyper}}=[y^{\mathrm{hyper}}_{1},\dots,y^{\mathrm{hyper}}_{C}]^{\top}$, the loss is computed as the sum of binary Focal Loss for each class:
$\mathrm{FL}(p,y)=-\alpha y(1-p)^{\gamma}\log(p)-(1-y)(1-\alpha)p^{\gamma}\log(1-p).$ (8)
@@
$\mathcal{L}_{i}=\delta_{i}\;\underbrace{\bigl[-\log h_{i,k_{i}}-\log S_{i,k_{i}}\bigr]}_{\mathcal{L}_{\text{event},i}}\;+\;(1-\delta_{i})(1-\alpha)\;\underbrace{\bigl[-\log S_{i,k_{i}^{\mathrm{c}}}\bigr]}_{\mathcal{L}_{\text{cens},i}},$ (10)
@@ D.2 Construction of Isolated Mask
[…]
@@
$\mathbf{A}_{p}\;=\;-\infty\;\Bigl[\mathbf{1}_{L\times L}\;-\;(\mathbf{m}_{p}\,\mathbf{m}_{p}^{\top})\,\odot\,\mathbf{E}_{p}\Bigr],\quad(\mathbf{E}_{p})_{ij}=\mathbb{I}\{(\mathbf{b}_{p})_{i}=(\mathbf{b}_{p})_{j}\}.$ (15)
@@ D.3 Dynamic Pack Length Adaptation
While the pack operation utilizes a fixed length $L$ for efficient batch processing, the actual number of instances sampled from each bag ($|\tilde{\mathcal{R}}_{b}|$ and $|\tilde{\mathcal{D}}_{b}|$) can vary significantly due to stochastic sampling and the diverse sizes of original WSIs. To accommodate this inherent variability, particularly when a mini-batch contains bags that yield a large number of sampled instances, we incorporate a dynamic pack length adaptation mechanism. Before packing the instances for a given branch (main or residual) within a mini-batch, we assess the maximum sampled sequence length from any single bag in that batch. Specifically, we check if $\max_{b}|\tilde{\mathcal{R}}_{b}|$ (for the main branch) or $\max_{b}|\tilde{\mathcal{D}}_{b}|$ (for the residual branch) exceeds the current pack length $L$. If this condition is met, the pack length for that specific branch and mini-batch is dynamically doubled to $2L$. This dynamic doubling occurs at most once per branch per mini-batch processing step, effectively setting an upper bound of $2L$ on the pack length. This adaptation ensures that sampled instances from bags with particularly large retained or discarded sets are less likely to be fragmented across numerous packs, leading to more efficient packing and potentially better representation within packs for such cases.
@@ E.1 Pack-based Batched Training
Training on variable-length sequences has traditionally relied on padding shorter sequences and applying masks to ignore padded tokens, ensuring uniform batch shapes at the cost of wasted computation Krell et al. (2021). To reduce this overhead, dynamic batching (length-based bucketing) groups sequences of similar lengths per batch, greatly minimizing padding requirements Żelasko et al. (2025). Modern transformer architectures further exploit attention masks to handle padding, and recent work goes beyond simple padding by packing multiple sequences into one longer sequence with special separators and adjusted position indices Krell et al. (2021). Such sequence packing techniques, originally used in large-scale NLP pre-training, can double throughput by eliminating pad tokens Kosec et al. (2021) while maintaining model fidelity via careful masking to prevent cross-sequence attention. For example, packing algorithms in BERT pre-training combine several short sentences into a single 512-token input, yielding  2× speedups with negligible accuracy loss Kosec et al. (2021). In computer vision, analogous ideas enable variable-resolution training. NaViT avoids fixed-size resizing by treating images as sequences of patches and packing arbitrary-resolution inputs, improving efficiency in large-scale image–text pre-training without sacrificing performance Dehghani et al. (2023). Other works dynamically reduce sequence length during processing, such as Token Merging (ToMe) merges redundant tokens in ViTs to halve the token count on the fly, boosting throughput  2× for large models with minimal accuracy drop Bolya et al. (2022). RNN-based systems commonly use packed sequences to skip computation on padded timesteps, and Transformer-based LLMs and vision models use padding masks or adaptive token pruning to similar effect. In CPath, where WSI yields a bag of thousands of instances, efficient batching is critical. However, the data characteristics in CPath render the direct application of the aforementioned strategies non-trivial. Approaches focused on packing short sequences provide limited benefits, while sampling long sequences risks significant information loss. Effectively adapting these efficient training paradigms to CPath is thus a key challenge. Our proposed pack-based framework addresses this by incorporating a residual branch to more effectively packing these variable-length long sequences, aiming to mitigate these limitations.
@@ E.2 More about Batchsize in Computational Pathology
As elaborated in the Related Work section, current slide-level MIL methods often train with batchsize of 1. Conversely, when explicit patch-level annotations are available for segmentation or detection tasks, researchers commonly employ moderate to large batch sizes by independently processing uniformly-sized patches extracted from WSIs, enabling stable training and efficient convergence Ciga et al. (2021), Graham et al. (2019). These models independently process uniformly-sized patches extracted from WSIs, leveraging moderate to large batch sizes to facilitate stable training and efficient convergence Ciga et al. (2021), Graham et al. (2019). Conversely, when explicit patch-level annotations are available for segmentation or detection tasks, researchers typically form batches at the patch level. They independently process uniformly-sized patches from WSIs using moderate to large batchsizes, facilitating stable training and efficient convergence Ciga et al. (2021), Graham et al. (2019). Such patch-centric batching, however, presents a discrepancy with the holistic slide assessment in clinical workflows.
Appendix F Limitation & Broader Impacts
@@
Appendix G Data Availability Statement

### v2

@@ B.7 Inference Time Comparison
B.8 Supplementary Confidence Intervals
@@
(a) Grading on PANDA & Sub-typing on BRCA
Method Grading (Acc.) Sub-typing (AUC)
CHIEF UNI GIGAP CHIEF UNI GIGAP
+RS 74.72±0.64 77.91±0.58 78.97±0.58 88.72±5.39 93.89±3.89 93.78±3.92
ABMIL +PackMIL 76.46±0.61 80.19±0.53 80.41±0.53 92.38±4.12 94.86±3.91 94.86±3.68
+RS 75.00±0.65 78.59±0.61 78.60±0.63 91.62±4.49 93.29±4.14 94.04±3.97
DSMIL +PackMIL 75.84±0.57 79.68±0.50 79.10±0.55 93.01±3.92 94.62±3.72 94.65±3.14
+RS 73.68±0.71 76.94±0.63 76.15±0.67 90.75±4.73 94.07±3.89 93.95±3.71
TransMIL +PackMIL 74.75±0.66 78.87±0.57 78.88±0.58 92.31±4.23 94.37±3.83 94.12±3.92
+RS 70.32±0.70 75.04±0.60 75.13±0.63 91.55±4.67 94.02±3.89 93.70±4.07
RRTMIL +PackMIL 74.63±0.69 78.46±0.59 78.43±0.60 92.43±3.92 94.54±3.79 94.47±4.01
(b) Survival Analysis on BRCA & LUAD
Method Survival-BRCA (C-index) Survival-LUAD (C-index)
CHIEF UNI GIGAP CHIEF UNI GIGAP
+RS 64.02±9.2 65.71±9.9 63.88±9.8 61.98±8.7 60.34±8.8 61.01±8.6
ABMIL +PackMIL 68.30±8.8 68.14±9.8 67.04±9.1 63.72±8.6 62.60±8.5 61.58±8.7
+RS 65.00±9.1 64.72±10.0 65.69±9.8 63.28±8.6 61.53±9.1 61.00±8.6
DSMIL +PackMIL 69.76±8.5 70.00±8.6 68.03±9.1 64.10±8.6 62.18±8.7 62.44±8.4
+RS 66.39±8.8 65.14±9.8 65.52±9.7 63.53±8.8 62.12±8.7 61.03±8.4
TransMIL +PackMIL 68.08±8.6 68.44±8.9 66.80±9.1 64.01±9.1 63.61±8.5 63.04±8.3
+RS 66.11±9.3 64.68±10.1 65.19±9.0 62.52±8.6 61.44±8.9 61.35±8.5
RRTMIL +PackMIL 68.15±9.1 68.73±9.4 67.62±9.0 64.37±8.4 62.01±8.6 62.79±8.2
@@ D.1 Task-specific Hyperslide Loss
Grading task. This task is modeled as single-label classification, we employ Asymmetric Loss (ASLLoss) [ridnik2021asymmetric]. This loss function is chosen to effectively address potential class imbalance and encourage the model to focus on correctly identifying positive classes while being less sensitive to negative misclassifications. Given the predicted probability distribution $\mathbf{p}=[p_{1},\dots,p_{G}]$ for $G$ classes and the corresponding one-hot encoded ground truth labels $\mathbf{y}^{\mathrm{hyper}}$, the loss is computed as:
$\begin{split}L_{\mathrm{grade}}(\mathbf{p},\mathbf{y}^{\mathrm{hyper}})=&-\sum_{i=1}^{G}\mathbf{y}^{\mathrm{hyper}}_{i}(1-p_{i})^{\gamma_{p}}\log(p_{i})\\ &-\sum_{i=1}^{G}(1-\mathbf{y}^{\mathrm{hyper}}_{i})p_{i}^{\gamma_{n}}\log(1-p_{i}),\end{split}$ (6)
@@
Subtyping task. This task is modeled as multi-label classification with soft targets, we use multi-label Focal Loss [lin2017focal]. This loss helps mitigate the issue of imbalanced subtype frequencies and focuses the model’s attention on harder-to-classify samples. Given the predicted probability vector $\mathbf{p}=[p_{1},\dots,p_{C}]^{\top}$ for $C$ subtypes (typically obtained via Sigmoid activation) and the corresponding soft ground truth label vector $\mathbf{y}^{\mathrm{hyper}}=[y^{\mathrm{hyper}}_{1},\dots,y^{\mathrm{hyper}}_{C}]^{\top}$, the loss is computed as the sum of binary Focal Loss for each class:
$\begin{split}\mathrm{FL}(p,y)=&-\alpha y(1-p)^{\gamma}\log(p)\\ &-(1-y)(1-\alpha)p^{\gamma}\log(1-p).\end{split}$ (8)
@@
$\begin{split}\mathcal{L}_{i}=&\delta_{i}\;\underbrace{\bigl[-\log h_{i,k_{i}}-\log S_{i,k_{i}}\bigr]}_{\mathcal{L}_{\text{event},i}}\;+\;\\ &(1-\delta_{i})(1-\alpha)\;\underbrace{\bigl[-\log S_{i,k_{i}^{\mathrm{c}}}\bigr]}_{\mathcal{L}_{\text{cens},i}},\end{split}$ (10)
@@ D.2 Construction of Isolated Mask
$\begin{split}(\mathbf{M}_{p})_{j,b}=(\mathbf{m}_{p})_{j}\cdot\mathbb{I}\{(\mathbf{b}_{p})_{j}=b\},\quad\\ \text{for }j=1,\dots,L,\;b=1,\dots,B.\end{split}$ (14)
@@
$\mathbf{A}_{p}\;=\;-\infty\;\Bigl[\mathbf{1}_{L\times L}\;-\;(\mathbf{m}_{p}\,\mathbf{m}_{p}^{\top})\,\odot\,\mathbf{E}_{p}\Bigr],$ (15)
$(\mathbf{E}_{p})_{ij}=\mathbb{I}\{(\mathbf{b}_{p})_{i}=(\mathbf{b}_{p})_{j}\}.$ (16)
@@ D.3 Dynamic Pack Length Adaptation
While the pack operation utilizes a fixed length $L$ for efficient batch processing, the actual number of instances sampled from each bag ($|\tilde{\mathcal{K}}_{b}|$ and $|\tilde{\mathcal{D}}_{b}|$) can vary significantly due to stochastic sampling and the diverse sizes of original WSIs. To accommodate this inherent variability, particularly when a mini-batch contains bags that yield a large number of sampled instances, we incorporate a dynamic pack length adaptation mechanism. Before packing the instances for a given branch (main or residual) within a mini-batch, we assess the maximum sampled sequence length from any single bag in that batch. Specifically, we check if $\max_{b}|\tilde{\mathcal{K}}_{b}|$ (for the main branch) or $\max_{b}|\tilde{\mathcal{D}}_{b}|$ (for the residual branch) exceeds the current pack length $L$. If this condition is met, the pack length for that specific branch and mini-batch is dynamically doubled to $2L$. This dynamic doubling occurs at most once per branch per mini-batch processing step, effectively setting an upper bound of $2L$ on the pack length. This adaptation ensures that sampled instances from bags with particularly large retained or discarded sets are less likely to be fragmented across numerous packs, leading to more efficient packing and potentially better representation within packs for such cases.
@@
D.4 Pseudo-code
Algorithm 1 outlines the PyTorch-style pseudocode for the ADS module. Subsequently, Algorithm 2 details the logic of our packing strategy.
# x: input instance features, shape $[B,N,D]$
# k: downsampling factor
# W_L, W_P: learnable linear projections
# attn: attention score
# 1. Compute attention-driven residual features
alpha = attn(x).softmax(dim=1)
x = x + x * alpha
# 2. Linear transformation
x = x @ W_L
# 3. Prepare for Grouping (Training vs Inference)
if training:
# Stochastic mode: Shuffle for diversity
x = shuffle(x, dim=1)
else:
# Deterministic mode: Keep order for interpretability
pool_mode = None
# 4. Instance Unshuffle and Pooling
B, N, D = x.shape
groups = x.view(B, N // k, k, D) # Reshape into groups of size k
if mode == ’max’:
x_pooled = groups.max(dim=2).values
elif mode == ’random’:
rand_idx = torch.randint(0, k, (B, N // k))
x_pooled = groups.gather(2, rand_idx)
else:
# Inference
x_pooled = x
# 5. Final projection
x_out = x_pooled @ W_P
Algorithm 1 PyTorch-style pseudocode for ADS (Training vs. Inference)
# batch: list of WSI feature sets, each shape $N_{i},D]$
# L: target fixed pack length
# r: keep ratio for main branch
# ADS: Attention-driven Downsampler module
buffer_main, buffer_res = [], []
# 1. Process each bag: Sample and Downsample
for x in batch:
N = x.shape[0]
# Stochastic instance-level sampling
perm = torch.randperm(N)
num_keep = int(N * r)
idx_keep, idx_disc = perm[:num_keep], perm[num_keep:]
# Apply ADS to obtain compact representations (ensure min/max length constraint)
x_keep = ADS(x[idx_keep]) # Shape: $n_{keep},D]$
x_disc = ADS(x[idx_disc]) # Shape: $n_{disc},D]$
buffer_main.append(x_keep)
buffer_res.append(x_disc)
# 2. Sequential Packing Operation (Next-Fit Strategy)
def run_packing(feature_list, L):
packs = []
current_pack = []
current_len = 0
for x in feature_list: # x: current bag features, Shape $[n,D]$
n = x.shape[0]
# Check if current bag fits in current pack
if current_len + n > L:
# Pack is full: Pad remaining slots with zeros
pad_size = L - current_len
finished_pack = concat(current_pack + [zeros(pad_size, D)]) # Shape: $[L,D]$
packs.append(finished_pack)
# Start new pack with current bag
current_pack = [x]
current_len = n
else:
# Append bag to current pack without splitting
current_pack.append(x)
current_len = current_len + n
# Handle the last remaining pack
if current_len > 0:
pad_size = L - current_len
last_pack = concat(current_pack + [zeros(pad_size, D)]) # Shape: $[L,D]$
packs.append(last_pack)
return stack(packs) # Final Output Shape: $B^{\prime},L,D]$
P_main = run_packing(buffer_main, L)
P_res = run_packing(buffer_res, L)
Algorithm 2 PyTorch-style pseudocode for Stochastic Sampling and Sequential Packing
@@ E.1 Pack-based Batched Training
Training on variable-length sequences has traditionally relied on padding shorter sequences and applying masks to ignore padded tokens, ensuring uniform batch shapes at the cost of wasted computation [krell2021efficient]. To reduce this overhead, dynamic batching (length-based bucketing) groups sequences of similar lengths per batch, greatly minimizing padding requirements [zelasko2025emmett]. Modern transformer architectures further exploit attention masks to handle padding, and recent work goes beyond simple padding by packing multiple sequences into one longer sequence with special separators and adjusted position indices [krell2021efficient]. Such sequence packing techniques, originally used in large-scale NLP pre-training, can double throughput by eliminating pad tokens [kosec2021packing] while maintaining model fidelity via careful masking to prevent cross-sequence attention. For example, packing algorithms in BERT pre-training combine several short sentences into a single 512-token input, yielding  2× speedups with negligible accuracy loss [kosec2021packing]. In computer vision, analogous ideas enable variable-resolution training. NaViT avoids fixed-size resizing by treating images as sequences of patches and packing arbitrary-resolution inputs, improving efficiency in large-scale image–text pre-training without sacrificing performance [dehghani2023patch]. Other works dynamically reduce sequence length during processing, such as Token Merging (ToMe) merges redundant tokens in ViTs to halve the token count on the fly, boosting throughput  2× for large models with minimal accuracy drop [bolya2022token]. RNN-based systems commonly use packed sequences to skip computation on padded timesteps, and Transformer-based LLMs and vision models use padding masks or adaptive token pruning to similar effect. In CPath, where WSI yields a bag of thousands of instances, efficient batching is critical. However, the data characteristics in CPath render the direct application of the aforementioned strategies non-trivial. Approaches focused on packing short sequences provide limited benefits, while sampling long sequences risks significant information loss. Effectively adapting these efficient training paradigms to CPath is thus a key challenge. Our proposed pack-based framework addresses this by incorporating a residual branch to more effectively packing these variable-length long sequences, aiming to mitigate these limitations.
@@ E.2 More about Batchsize in Computational Pathology
As elaborated in the Related Work section, current slide-level MIL methods often train with batchsize of 1. Conversely, when explicit patch-level annotations are available for segmentation or detection tasks, researchers commonly employ moderate to large batch sizes by independently processing uniformly-sized patches extracted from WSIs, enabling stable training and efficient convergence [ciga2021overcoming, graham2019hover]. These models independently process uniformly-sized patches extracted from WSIs, leveraging moderate to large batch sizes to facilitate stable training and efficient convergence [ciga2021overcoming, graham2019hover]. Conversely, when explicit patch-level annotations are available for segmentation or detection tasks, researchers typically form batches at the patch level. They independently process uniformly-sized patches from WSIs using moderate to large batchsizes, facilitating stable training and efficient convergence [ciga2021overcoming, graham2019hover]. Such patch-centric batching, however, presents a discrepancy with the holistic slide assessment in clinical workflows.
Appendix F Conclusion
Appendix G Limitation & Broader Impacts
@@
Appendix H Data Availability Statement